# Rank Two Sheaves With Low Discriminant on the Fano Threefold of Index 2 and Degree 5

### v1

@@ 3. Objects with Maximal Third Chern Character
Suppose that $E$ is destabilized at a semicircular wall by a tilt-stable subobject or quotient $F$ with $\mathrm{ch}_{\leq 2}(F)=s+xH+yH^{2}$. The wall should intersect the ray $\beta=\beta_{-}(E)=-\frac{1}{2}-\sqrt{\frac{3}{20}}$. Again we may assume that $s>0,\mu(F)<\mu(E)$. Since $0<\mathrm{ch}_{1}^{\beta_{-}(E)}(F)<\mathrm{ch}_{1}^{\beta_{-}(E)}(E)$, we have
(10) $0<x+\frac{s}{2}+\sqrt{\frac{3}{20}}s<\sqrt{\frac{3}{5}}.$
@@
Tilt-stability of $E$ implies that $H^{0}(E)=0$. If $H^{2}(E)\neq 0$, then there would exist a non-zero morphism $E\to\mathcal{O}_{X}(-2)[1]$, contradicting the fact that $E$ is tilt-stable for all $\beta<-\frac{1}{2},\alpha>0$. Grothendieck-Riemann-Roch Theorem gives that $\chi(E)=5e-\frac{1}{6}\leq 0$. This gives $e\leq\frac{1}{30}$. In case of equality Lemma 2.10 and Corollary 2.15 imply that $E[1]$ belongs to $\mathfrak{C}$. Chern character of $E$ implies that $E\cong\mathcal{U}$. ∎
Lemma 3.7.
If $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable in $\overline{U}$ for a non-empty open subset $U\subset\mathbb{H}$ and $\mathrm{ch}(E)=2-\frac{H^{2}}{5}+eH^{3}$, then $e\leq-\frac{1}{5}$, and in case of equality $E$ is a coherent sheaf, included into an exact triple
(11) $0\to E\to\mathcal{O}_{X}^{\oplus 2}\to\mathcal{O}_{L}(1)\to 0$
@@
Proof.
In this case $\mathrm{ch}_{\leq 2}(E)$ also cannot be an integer multiple of $\mathrm{ch}_{\leq 2}$ of another object, hence $E$ is tilt-stable in $U$. Any semicircular wall for $E$ has to intersect the ray $\beta=\beta_{-}(E)=-\frac{\sqrt{5}}{5}$. We can assume that this wall is given by a tilt-stable subobject or quotient $F$ with $\mathrm{ch}_{\leq 2}(F)=s+xH+yH^{2}$, $s>0,\mu(F)<\mu(E)$. Since $0<\mathrm{ch}_{1}^{\beta_{-}(E)}(F)<\mathrm{ch}_{1}^{\beta_{-}(E)}(E)$, we have
(12) $0<x+\frac{\sqrt{5}}{5}s<\frac{2\sqrt{5}}{5}.$
@@
If $s=1$, then (12) implies that $\frac{\sqrt{5}}{5}<x<\frac{\sqrt{5}}{5}$, so $x=0$, but in this case the property $\mu(F)<\mu(E)$ does not hold and the wall is not semicircular. If $s=2$, then by (12) 1<x<0-1<x<0, which is impossible. And if $s=3$, then (12) implies that 2<x<0-2<x<0, so x=1x=-1. Inequalities $0\leq\overline{\Delta}_{H}(F)<20$ together with the fact that $y\in-\frac{1}{2}+\frac{1}{5}\mathbb{Z}$ (since $c_{i}(F)\in H^{2i}(X,\mathbb{Z})$) imply that $y=\frac{1}{10}$. But then we find that $\overline{\Delta}_{H}(F)=\overline{\Delta}_{H}(E/F)=10$, contradicting the inequality $\overline{\Delta}_{H}(F)+\overline{\Delta}_{H}(E/F)<\overline{\Delta}_{H}(E)$. Therefore, $E$ is $\nu_{\alpha,\beta}$-stable for all $\alpha>0,\beta<0$.
Tilt-stability of $E$ implies that $H^{0}(E)=0$. Also $H^{2}(E)\cong\operatorname{Hom}(E,\mathcal{O}_{X}(-2)[1])^{\vee}=0$, because otherwise there would be a semicircular wall. Grothendieck-Riemann-Roch Theorem gives $\chi(E)=5e+1\leq 0$, hence $e\leq-\frac{1}{5}$. Assume that $e=-\frac{1}{5}$. Proposition 2.7 implies that $E$ is a 2-stable sheaf. Since $E$ is a rank 2 sheaf with negative third Chern class, it cannot be reflexive (see the proof of [11, Proposition 2.6]). The sheaf $E$ can be included into an exact sequence
@@
where $\dim\mathrm{supp}(T)\leq 1$. If $\dim\mathrm{supp}(T)=0$, then $\nu_{\alpha,\beta}(E^{\vee\vee})=\nu_{\alpha,\beta}(E)$ for all $(\beta,\alpha)$, and tilt-stability of $E$ implies tilt-stability of $E^{\vee\vee}$. Applying to $E^{\vee\vee}$ the above arguments, we obtain that $\frac{\mathrm{ch}_{3}(E^{\vee\vee})}{H^{3}}\leq-\frac{1}{5}$, contradicting reflexivity of $E^{\vee\vee}$.
@@
Since $\mathrm{ch}(T)=\frac{H^{2}}{5}+\frac{H^{3}}{5}$, one-dimensional part of $\mathrm{supp}(T)$ consists of one projective line $L\subset X$. Classification of coherent sheaves on $\mathbb{P}^{1}$ implies that $T\cong\mathcal{O}_{L}(a)\oplus T^{\prime},\dim\mathrm{supp}(T^{\prime})\leq 0$. Chern character of $T$ together with existence of a surjective morphism $\mathcal{O}_{X}^{\oplus 2}\twoheadrightarrow T$ implies that $T\cong\mathcal{O}_{L}(1)$ or $T\cong\mathcal{O}_{L}\oplus\mathcal{O}_{p},p\in X$. However, in the second case one can, using that $H^{0}(E)=0$, construct a subsheaf $\mathcal{I}_{p}\hookrightarrow E$, contradicting 2-stability of $E$. And in case when $T\cong\mathcal{O}_{L}(1)$ one can show that $\operatorname{Hom}(\mathcal{I}_{Z},E)=0$ if $\dim Z\leq 0$, implying that $E$ is 2-stable. ∎
Lemma 3.8.
@@
If $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable in $\overline{U}$ for a non-empty open subset $U\subset\mathbb{H}$ and $\mathrm{ch}(E)=2-H-\frac{H^{2}}{10}+eH^{3}$, then $e\leq\frac{7}{30}$, and in case of equality $E$ is a coherent sheaf, included into an exact triple
(14) $0\to\mathcal{U}(-1)\to\mathcal{O}_{X}(-1)^{\oplus 4}\to E\to 0.$
Proof.
@@
Note that $H^{2}\cdot\mathrm{ch}^{-1}_{1}(E)$ has the least positive value, so there is no wall for $E$ at $\beta=-1$ for any $\alpha>0$. The inequality $Q_{\alpha,\beta}(E)<0$ defines a semi-disk lying below the numerical wall $Q_{\alpha,\beta}(E)=0$. Therefore the generalized Bogomolov-Gieseker inequality together with the fact that the walls are nested implies that $Q_{0,-1}(E)=41-150e\geq 0$. Using also the integrality of Chern classes of $E$, we obtain the bound $e\leq\frac{7}{30}$. Assume that $e=\frac{7}{30}$.
@@
A calculation shows that the numerical wall $W(E,\mathcal{O}_{X}(-3)[1])$ intersects the ray $\beta=-1$. Therefore, $\mathrm{Ext}^{2}(\mathcal{O}_{X}(-1),E)\cong\operatorname{Hom}(E,\mathcal{O}_{X}(-3)[1])=0$. Grothendieck-Riemann-Roch Theorem gives that $\operatorname{Hom}(\mathcal{O}_{X}(-1),E)\geq\chi(\mathcal{O}_{X}(-1),E)=4$. This gives an exact sequence
(15) $0\to\mathcal{O}_{X}(-1)^{\oplus 4}\to E\to G\to 0$
@@
in the category $\mathrm{Coh}^{\beta}(X)$ for $\beta<-1$ large enough. Here $\mathrm{ch}(G)=-2+3H-\frac{21}{10}H^{2}+\frac{9}{10}H^{3}$. We find $\mathrm{ch}(\mathbf{R}\mathcal{H}om(G,\mathcal{O}_{X})[1](-2))=2-H+\frac{H^{2}}{10}+\frac{H^{3}}{30}$. Proposition 3.2 gives an exact sequence
$0\to\widetilde{G}\to\mathbf{R}\mathcal{H}om(G,\mathcal{O}_{X})[1](-2)\to T[-1]\to 0$
@@
with a tilt-semistable object $\widetilde{G}$ and a sheaf $T$ with $\dim T\leq 0$. It follows that $\frac{\mathrm{ch}_{3}(\widetilde{G})}{H^{3}}=\frac{1}{30}+\frac{\mathrm{ch}_{3}(T)}{H^{3}}.$ Since $\frac{\mathrm{ch}_{3}(T)}{H^{3}}\geq 0$ and, by Lemma 3.6, $\frac{\mathrm{ch}_{3}(\widetilde{G})}{H^{3}}\leq\frac{1}{30}$, we obtain that $T=0$ and $\mathbf{R}\mathcal{H}om(G,\mathcal{O}_{X})[1](-2)\cong\mathcal{U}$. It follows that $G\cong\mathcal{U}(-1)[1]$. A calculation shows that $\nu_{\alpha,\beta}(\mathcal{O}_{X}(-1))>\nu_{\alpha,\beta}(E)$ under the wall $W(\mathcal{O}_{X}(-1),E)$, hence $E$ is not tilt-semistable there. Since any wall lying above $W(\mathcal{O}_{X}(-1),E)$ would intersect the ray $\beta=-1$, it follows that $E$ is tilt-semistable everywhere above $W(\mathcal{O}_{X}(-1),E)$. By Proposition 2.7 $E$ is a sheaf, and (15) gives the exact triple (14). ∎
@@
Lemma 3.9.
If $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable in $\overline{U}$ for a non-empty open subset $U\subset\mathbb{H}$ and $\mathrm{ch}(E)=3-H+\frac{1}{10}H^{2}+eH^{3}$, then $e\leq-\frac{1}{6}$.
Proof.
Any semicircular wall for $E$ has to intersect the ray $\beta=\beta_{-}(E)=\frac{-10-2\sqrt{10}}{30}$. Assume that this wall is given by a tilt-stable subobject or quotient $F$ with $\mathrm{ch}_{\leq 2}(F)=s+xH+yH^{2},s>0,\mu(F)<\mu(E)$. Since $0<\mathrm{ch}_{1}^{\beta_{-}(E)}(F)<\mathrm{ch}_{1}^{\beta_{-}(E)}(E)$, we have
(16) $0<x+\frac{1}{3}s+\frac{2\sqrt{10}}{30}s<\frac{2\sqrt{10}}{10}.$
@@
We proved that there are no semicircular walls for $E$. 2-stability of $E$ implies that $H^{0}(E)=0$, and the absense of semicircular walls implies $H^{2}(E)=\operatorname{Hom}(E,\mathcal{O}_{X}(-2)[1])^{\vee}=0$. A calculation gives $\chi(E)=5e+\frac{5}{6}\leq 0$, hence $e\leq-\frac{1}{6}$. ∎
Lemma 3.10.
If $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable in $\overline{U}$ for a non-empty open subset $U\subset\mathbb{H}$ and $\mathrm{ch}(E)=3+H+\frac{1}{10}H^{2}+eH^{3}$, then $e\leq-\frac{2}{15}$.
Proof.
We have $\mathrm{ch}(E(-1))=3-2H+\frac{3}{5}H^{2}+(e-\frac{1}{10})H^{3}$, and E(1)E(-1) also has to be tilt-semistable in a subset of $\mathbb{H}$ with non-empty interior. So Lemma 3.5 gives $e\leq-\frac{2}{15}$. ∎
Corollary 3.11.
There are no objects $E\in\mathrm{Coh}^{\beta}(X)$ with $\mathrm{ch}_{\leq 2}(E)=3-H+\frac{1}{10}H^{2}$, which are tilt-semistable in $\overline{U}$ for some non-empty open subset $U\subset\mathbb{H}$.
Proof.
By the proof of Lemma 3.9 such an object $E$ should be tilt-stable for all $\beta<-\frac{1}{3},\alpha>0$, hence $E$ is a 2-stable sheaf. The dual sheaf $E^{\vee}$ should be slope semistable, and in fact slope stable since its rank and first Chern class are coprime. If $E$ is reflexive, then $\mathrm{ch}_{\leq 2}(E^{\vee})=3+H+\frac{1}{10}H^{2}$. However, in this case Lemmas 3.9 and 3.10 imply that $\mathrm{ch}_{3}(E)+\mathrm{ch}_{3}(E^{\vee})<0$, contradicting Lemma 2.17, so the sheaf $E$ is not reflexive. There is a short exact sequence
@@
with $\dim\mathrm{supp}\ T\leq 1$. The support of $T$ cannot be zero-dimensional, because in this case $\mathrm{ch}_{\leq 2}(E^{\vee\vee})=3-H+\frac{1}{10}H^{2}$ and by the previous argument $E^{\vee\vee}$ cannot be reflexive. And if $\dim\mathrm{supp}\ T=1$, then $\mathrm{ch}_{\leq 2}(E^{\vee\vee})=3-H+(\frac{1}{10}+d)H^{2}$ for $d\geq\frac{1}{5}$ and Bogomolov-Gieseker inequality for $E^{\vee\vee}$ is not satisfied. ∎
@@
Lemma 3.12.
If $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable in $\overline{U}$ for a non-empty open subset $U\subset\mathbb{H}$ and $\mathrm{ch}(E)=2-\frac{2}{5}H^{2}+eH^{3}$, then $e\leq 0$, and in case of equality $E$ is a coherent sheaf, included into an exact triple
(17) $0\to\mathcal{Q}(-1)^{\oplus 2}\to\mathcal{U}^{\oplus 4}\to E\to 0.$
Proof.
Assume, firstly, that $E$ is tilt-stable in $U$. Note that $\beta_{-}(E)=-\frac{\sqrt{10}}{5}\in(-1,-\frac{1}{2})$. In order to use the heart $\mathfrak{C}$, we need to find a point $(\beta_{0},\alpha_{0})$ on the hyperbola $\nu_{\alpha,\beta}(E)=-\frac{2}{5}+\beta^{2}-\alpha^{2}=0$ with $\alpha_{0}<\beta_{0}+1$, such that $E$ is $\nu_{{\alpha_{0}},{\beta_{0}}}$-semistable. The curve $\frac{2}{5}+\beta^{2}-\alpha^{2}=0$ intersects the line $\alpha=\beta+1$ at the point with $\beta=-\frac{7}{10},\alpha=\frac{3}{10}$, hence it suffices to find such a point $(\beta_{0},\alpha_{0})$ with $\alpha_{0}<\frac{3}{10}$.
@@
(18) $0<x+\frac{\sqrt{10}}{5}s<\frac{2\sqrt{10}}{5}.$
@@
Tilt-semistability of $E$ implies that $H^{0}(E)=0$. If $H^{2}(E)\neq 0$, then there is a nonzero morphism $E\to\mathcal{O}_{X}(-2)[1]$. However, such a morphism would destabilize $E$ everywhere below the wall $W(E,\mathcal{O}_{X}(-2)[1])$, which is located above $W(E,\mathcal{O}_{X}(-1)[1])$ and has radius bigger than $\frac{3}{10}$, contradicting what we have proved. A calculation by Grothendieck-Riemann-Roch Theorem gives $\chi(E)=5e\leq 0$, hence $e\leq 0$. Since for $e=0$ $\mathrm{ch}(E[1])=4\mathrm{ch}(\mathcal{U}[1])+2\mathrm{ch}(\mathcal{Q}(-1)[2])$, we get an exact sequence of the form (17).
Suppose now that $E$ is strictly tilt-semistable in $U$. In this case $E$ is an extension
$0\to F\to E\to G\to 0$
with $\mathrm{ch}_{\leq 2}(F)=\mathrm{ch}_{\leq 2}(G)=1-\frac{H^{2}}{5}$, the objects $F$ and $G$ are tilt-stable in $U$. The curve $\nu_{\alpha,\beta}(F)=0$ is the same hyperbola $\frac{2}{5}+\beta^{2}-\alpha^{2}=0$ as in the case of $E$. Note that $\mathrm{ch}_{1}^{-\frac{1}{2}}(F)$ has the least positive value, hence there are no semicircular walls for $F$ intersecting the ray $\beta=-\frac{1}{2}$. The numerical wall $W(F,\mathcal{O}_{X}(-1))$ intersects this ray, hence any wall for $F$ should lie below $W(F,\mathcal{O}_{X}(-1))$ and its radius has to be smaller than $\frac{3}{10}$. The same arguments as above imply that $\mathrm{ch}_{3}(F)\leq 0,F[1]\in\mathfrak{C}$, analogously $\mathrm{ch}_{3}(G)\leq 0,G[1]\in\mathfrak{C}$, hence in this case $E[1]$ also belongs to $\mathfrak{C}$, and in the case of maximal $\mathrm{ch}_{3}(E)$ we get the exact triple (17). ∎
Remark 3.13.
@@
Using the Serre correspondence, one can prove that there exist slope stable rank 2 vector bundles $E$ on $X$ with $c_{1}(E)=0,c_{2}(E)=2$ and $H^{1}(E(-1))=0$, the so-called instanton bundles [9, Theorem D, Step 1 of the proof]. It follows that there indeed exist tilt-stable sheaves $E$ from (17).
@@
Proof of Theorem 3.4. The bounds on $\mathrm{ch}_{2}(E)$ follow from Bogomolov-Gieseker Inequality. Other statements are proven in Proposition 3.1 and Lemmas 3.6, 3.7, 3.8, 3.12.∎
4. Partial Results and a Conjecture
Lemma 4.1.

### v2

@@ 3 Objects with Maximal Third Chern Character
Proof. Suppose that $E$ is destabilized at a semicircular wall by a tilt-stable subobject or quotient $F$ with $\mathrm{ch}_{\leq 2}(F)=s+xH+yH^{2}$. The wall should intersect the ray $\beta=\beta_{-}(E)=-\frac{1}{2}-\sqrt{\frac{3}{20}}$. Again we may assume that $s>0,\mu(F)<\mu(E)$. Since $0<\mathrm{ch}_{1}^{\beta_{-}(E)}(F)<\mathrm{ch}_{1}^{\beta_{-}(E)}(E)$, we have
$0<x+\frac{s}{2}+\sqrt{\frac{3}{20}}s<\sqrt{\frac{3}{5}}.$ (11)
@@
Tilt-stability of $E$ implies that $H^{0}(E)=0$. If $H^{2}(E)\neq 0$, then there would exist a non-zero morphism $E\to\mathcal{O}_{X}(-2)[1]$, contradicting the fact that $E$ is tilt-stable for all $\beta<-\frac{1}{2},\alpha>0$. Grothendieck-Riemann-Roch Theorem gives that $\chi(E)=5e-\frac{1}{6}\leq 0$. This gives $e\leq\frac{1}{30}$. In case of equality Lemma 2.10 and Corollary 2.15 imply that $E[1]$ belongs to $\mathfrak{C}$. Chern character of $E$ implies that $E\cong\mathcal{U}$. $\square$
Lemma 3.7
If $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable and $\mathrm{ch}(E)=2-\frac{H^{2}}{5}+eH^{3}$, then $e\leq-\frac{1}{5}$, and in case of equality $E$ is a coherent sheaf, included into an exact triple
$0\to E\to\mathcal{O}_{X}^{\oplus 2}\to\mathcal{O}_{L}(1)\to 0$ (12)
@@
Proof. Any semicircular wall for $E$ has to intersect the ray $\beta=\beta_{-}(E)=-\frac{\sqrt{5}}{5}$. We can assume that this wall is given by a tilt-stable subobject or quotient $F$ with $\mathrm{ch}_{\leq 2}(F)=s+xH+yH^{2}$, $s>0,\mu(F)<\mu(E)$. Since $0<\mathrm{ch}_{1}^{\beta_{-}(E)}(F)<\mathrm{ch}_{1}^{\beta_{-}(E)}(E)$, we have
$0<x+\frac{\sqrt{5}}{5}s<\frac{2\sqrt{5}}{5}.$ (13)
@@
If $s=1$, then (13) implies that $\frac{\sqrt{5}}{5}<x<\frac{\sqrt{5}}{5}$, so $x=0$, but in this case the property $\mu(F)<\mu(E)$ does not hold and the wall is not semicircular. If $s=2$, then by (13) 1<x<0-1<x<0, which is impossible. And if $s=3$, then (13) implies that 2<x<0-2<x<0, so x=1x=-1. Inequalities $0\leq\overline{\Delta}_{H}(F)<20$ together with the fact that $y\in-\frac{1}{2}+\frac{1}{5}\mathbb{Z}$ (since $c_{i}(F)\in H^{2i}(X,\mathbb{Z})$) imply that $y=\frac{1}{10}$. But then we find that $\overline{\Delta}_{H}(F)=\overline{\Delta}_{H}(E/F)=10$, contradicting the inequality $\overline{\Delta}_{H}(F)+\overline{\Delta}_{H}(E/F)<\overline{\Delta}_{H}(E)$. Therefore, $E$ is $\nu_{\alpha,\beta}$-semistable for all $\alpha>0,\beta<0$.
Tilt-semistability of $E$ implies that $H^{0}(E)=0$. Also $H^{2}(E)\cong\mathrm{Hom}(E,\mathcal{O}_{X}(-2)[1])^{\vee}=0$, because otherwise there would be a semicircular wall. Grothendieck-Riemann-Roch Theorem gives $\chi(E)=5e+1\leq 0$, hence $e\leq-\frac{1}{5}$. Assume that $e=-\frac{1}{5}$. Proposition 2.7 implies that $E$ is a 2-semistable sheaf. Since $E$ is a rank 2 sheaf with negative third Chern class, it cannot be reflexive (see the proof of [10, Proposition 2.6]). The sheaf $E$ can be included into an exact sequence
@@
where $\dim\mathrm{supp}(T)\leq 1$. If $\dim\mathrm{supp}(T)=0$, then $\nu_{\alpha,\beta}(E^{\vee\vee})=\nu_{\alpha,\beta}(E)$ for all $(\beta,\alpha)$, and tilt-semistability of $E$ implies tilt-semistability of $E^{\vee\vee}$. Applying to $E^{\vee\vee}$ the above arguments, we obtain that $\frac{\mathrm{ch}_{3}(E^{\vee\vee})}{H^{3}}\leq-\frac{1}{5}$, contradicting reflexivity of $E^{\vee\vee}$.
@@
Since $\mathrm{ch}(T)=\frac{H^{2}}{5}+\frac{H^{3}}{5}$, one-dimensional part of $\mathrm{supp}(T)$ consists of one projective line $L\subset X$. Classification of coherent sheaves on $\mathbb{P}^{1}$ implies that $T\cong\mathcal{O}_{L}(a)\oplus T^{\prime},\dim\mathrm{supp}(T^{\prime})\leq 0$. Chern character of $T$ together with existence of a surjective morphism $\mathcal{O}_{X}^{\oplus 2}\twoheadrightarrow T$ implies that $T\cong\mathcal{O}_{L}(1)$ or $T\cong\mathcal{O}_{L}\oplus\mathcal{O}_{p},p\in X$. However, in the second case one can, using that $H^{0}(E)=0$, construct a subsheaf $\mathcal{I}_{p}\hookrightarrow E$, contradicting 2-stability of $E$. And in case when $T\cong\mathcal{O}_{L}(1)$ one can show that $\mathrm{Hom}(\mathcal{I}_{Z},E)=0$ if $\dim Z\leq 0$, implying that $E$ is 2-semistable. $\square$
Lemma 3.8
@@
If $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable and $\mathrm{ch}(E)=2-H-\frac{H^{2}}{10}+eH^{3}$, then $e\leq\frac{7}{30}$, and in case of equality $E$ is included into an exact triple
$0\to\mathcal{O}_{X}(-1)^{\oplus 4}\to E\to\mathcal{U}(-1)[1]\to 0.$ (15)
@@
Proof. Note that $H^{2}\cdot\mathrm{ch}^{-1}_{1}(E)$ has the least positive value, so there is no wall for $E$ at $\beta=-1$ for any $\alpha>0$. The inequality $Q_{\alpha,\beta}(E)<0$ defines a semi-disk lying below the numerical wall $Q_{\alpha,\beta}(E)=0$. Therefore the generalized Bogomolov-Gieseker inequality together with the fact that the walls are nested implies that $Q_{0,-1}(E)=41-150e\geq 0$. Using also the integrality of Chern classes of $E$, we obtain the bound $e\leq\frac{7}{30}$. Assume that $e=\frac{7}{30}$.
@@
A calculation shows that the numerical wall $W(E,\mathcal{O}_{X}(-3)[1])$ intersects the ray $\beta=-1$. Therefore, $\mathrm{Ext}^{2}(\mathcal{O}_{X}(-1),E)\cong\mathrm{Hom}(E,\mathcal{O}_{X}(-3)[1])=0$. Grothendieck-Riemann-Roch Theorem gives that $\mathrm{Hom}(\mathcal{O}_{X}(-1),E)\geq\chi(\mathcal{O}_{X}(-1),E)=4$. This gives an exact sequence
$0\to\mathcal{O}_{X}(-1)^{\oplus 4}\to E\to G\to 0$ (16)
@@
in the category $\mathrm{Coh}^{\beta}(X)$ for $\beta<-1$ large enough. Here $\mathrm{ch}(G)=-2+3H-\frac{21}{10}H^{2}+\frac{9}{10}H^{3}$. We find $\mathrm{ch}(\mathbf{R}\mathcal{H}om(G,\mathcal{O}_{X})[1](-2))=2-H+\frac{H^{2}}{10}+\frac{H^{3}}{30}$. Proposition 3.2 gives an exact triangle
$\widetilde{G}\to\mathbf{R}\mathcal{H}om(G,\mathcal{O}_{X})[1](-2)\to T[-1]\to\widetilde{G}[1]$
@@
with a tilt-semistable object $\widetilde{G}$ and a sheaf $T$ with $\dim T\leq 0$. It follows that $\frac{\mathrm{ch}_{3}(\widetilde{G})}{H^{3}}=\frac{1}{30}+\frac{\mathrm{ch}_{3}(T)}{H^{3}}.$ Since $\frac{\mathrm{ch}_{3}(T)}{H^{3}}\geq 0$ and, by Lemma 3.6, $\frac{\mathrm{ch}_{3}(\widetilde{G})}{H^{3}}\leq\frac{1}{30}$, we obtain that $T=0$ and $\mathbf{R}\mathcal{H}om(G,\mathcal{O}_{X})[1](-2)\cong\mathcal{U}$. It follows that $G\cong\mathcal{U}(-1)[1]$. $\square$
Remark 3.9
Note that the exact sequence (15) gives an exact triangle
$\mathcal{O}_{X}(-1)^{\oplus 4}\to E\to\mathcal{U}(-1)[1]\to\mathcal{O}_{X}(-1)^{\oplus 4}[1].$
Rotating it, we obtain an exact triangle
$\mathcal{U}(-1)\to\mathcal{O}_{X}(-1)^{\oplus 4}\to E\to\mathcal{U}(-1)[1].$
Hence, if $E$ is a sheaf, we have an exact sequence
$0\to\mathcal{U}(-1)\to\mathcal{O}_{X}(-1)^{\oplus 4}\to E\to 0.$
@@
Lemma 3.10
If $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable and $\mathrm{ch}(E)=3-H+\frac{1}{10}H^{2}+eH^{3}$, then $e\leq-\frac{1}{6}$.
Proof. Any semicircular wall for $E$ has to intersect the ray $\beta=\beta_{-}(E)=\frac{-10-2\sqrt{10}}{30}$. Assume that this wall is given by a tilt-stable subobject or quotient $F$ with $\mathrm{ch}_{\leq 2}(F)=s+xH+yH^{2},s>0,\mu(F)<\mu(E)$. Since $0<\mathrm{ch}_{1}^{\beta_{-}(E)}(F)<\mathrm{ch}_{1}^{\beta_{-}(E)}(E)$, we have
$0<x+\frac{1}{3}s+\frac{2\sqrt{10}}{30}s<\frac{2\sqrt{10}}{10}.$ (17)
@@
We proved that there are no semicircular walls for $E$. 2-stability of $E$ implies that $H^{0}(E)=0$, and the absense of semicircular walls implies $H^{2}(E)=\mathrm{Hom}(E,\mathcal{O}_{X}(-2)[1])^{\vee}=0$. A calculation gives $\chi(E)=5e+\frac{5}{6}\leq 0$, hence $e\leq-\frac{1}{6}$. $\square$
Lemma 3.11
If $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable and $\mathrm{ch}(E)=3+H+\frac{1}{10}H^{2}+eH^{3}$, then $e\leq-\frac{2}{15}$.
Proof. We have $\mathrm{ch}(E(-1))=3-2H+\frac{3}{5}H^{2}+(e-\frac{1}{10})H^{3}$, and E(1)E(-1) is also tilt-semistable. So Lemma 3.5 gives $e\leq-\frac{2}{15}$. $\square$
Corollary 3.12
There are no tilt-semistable objects $E\in\mathrm{Coh}^{\beta}(X)$ with $\mathrm{ch}_{\leq 2}(E)=3-H+\frac{1}{10}H^{2}$.
Proof. By the proof of Lemma 3.10 such an object $E$ should be tilt-semistable for all $\beta<-\frac{1}{3},\alpha>0$, hence $E$ is a 2-semistable sheaf. The dual sheaf $E^{\vee}$ should be slope semistable, and in fact slope stable since its rank and first Chern class are coprime. If $E$ is reflexive, then $\mathrm{ch}_{\leq 2}(E^{\vee})=3+H+\frac{1}{10}H^{2}$. However, in this case Lemmas 3.10 and 3.11 imply that $\mathrm{ch}_{3}(E)+\mathrm{ch}_{3}(E^{\vee})<0$, contradicting Lemma 2.17, so the sheaf $E$ is not reflexive. There is a short exact sequence
@@
with $\dim\mathrm{supp}\ T\leq 1$. The support of $T$ cannot be zero-dimensional, because in this case $\mathrm{ch}_{\leq 2}(E^{\vee\vee})=3-H+\frac{1}{10}H^{2}$ and by the previous argument $E^{\vee\vee}$ cannot be reflexive. And if $\dim\mathrm{supp}\ T=1$, then $\mathrm{ch}_{\leq 2}(E^{\vee\vee})=3-H+(\frac{1}{10}+d)H^{2}$ for $d\geq\frac{1}{5}$ and Bogomolov-Gieseker inequality for $E^{\vee\vee}$ is not satisfied. $\square$
@@
Lemma 3.13
If $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable and $\mathrm{ch}(E)=2-\frac{2}{5}H^{2}+eH^{3}$, then $e\leq 0$, and in case of equality $E$ is included into an exact triple
$0\to\mathcal{U}^{\oplus 4}\to E\to\mathcal{Q}(-1)^{\oplus 2}[1]\to 0.$ (18)
Proof. Note that $\beta_{-}(E)=-\frac{\sqrt{10}}{5}\in(-1,-\frac{1}{2})$. In order to use the heart $\mathfrak{C}$, we need to find a point $(\beta_{0},\alpha_{0})$ on the hyperbola $\nu_{\alpha,\beta}(E)=-\frac{2}{5}+\beta^{2}-\alpha^{2}=0$ with $\alpha_{0}<\beta_{0}+1$, such that $E$ is $\nu_{{\alpha_{0}},{\beta_{0}}}$-semistable. The curve $\frac{2}{5}+\beta^{2}-\alpha^{2}=0$ intersects the line $\alpha=\beta+1$ at the point with $\beta=-\frac{7}{10},\alpha=\frac{3}{10}$, hence it suffices to find such a point $(\beta_{0},\alpha_{0})$ with $\alpha_{0}<\frac{3}{10}$.
@@
$0<x+\frac{\sqrt{10}}{5}s<\frac{2\sqrt{10}}{5}.$ (19)
@@
Tilt-semistability of $E$ implies that $H^{0}(E)=0$. If $H^{2}(E)\neq 0$, then there is a nonzero morphism $E\to\mathcal{O}_{X}(-2)[1]$. However, such a morphism would destabilize $E$ everywhere below the wall $W(E,\mathcal{O}_{X}(-2)[1])$, which is located above $W(E,\mathcal{O}_{X}(-1)[1])$ and has radius bigger than $\frac{3}{10}$, contradicting what we have proved. A calculation by Grothendieck-Riemann-Roch Theorem gives $\chi(E)=5e\leq 0$, hence $e\leq 0$. Since for $e=0$ $\mathrm{ch}(E[1])=4\mathrm{ch}(\mathcal{U}[1])+2\mathrm{ch}(\mathcal{Q}(-1)[2])$, we get an exact sequence of the form (18) (using Remark 2.16). $\square$
Remark 3.14
@@
Using the Serre correspondence, one can prove that there exist slope stable rank 2 vector bundles $E$ on $X$ with $c_{1}(E)=0,c_{2}(E)=2$ and $H^{1}(E(-1))=0$, the so-called instanton bundles [8, Theorem D, Step 1 of the proof]. It follows that there exist tilt-stable sheaves $E$ from (18).
@@
Proof of Theorem 3.4. The bounds on $\mathrm{ch}_{2}(E)$ follow from Bogomolov-Gieseker Inequality. Other statements are proven in Proposition 3.1 and Lemmas 3.6, 3.7, 3.8, 3.13. $\square$
4 Partial Results and a Conjecture
Lemma 4.1